OpenStudy (anonymous):

PR has an endpoint at (25, -5) and a midpoint of (18, -1). What is the value of the x-coordinate of the other endpoint?

OpenStudy (anonymous):

These confuse me sooooo much

OpenStudy (raden):

let the other point is (x,y) (25+x)/2 = 18, solve for x also (-5+y)/2 = -1, solve for y

OpenStudy (anonymous):

x= 21.5? y= -2

OpenStudy (raden):

(25+x)/2 = 18 multiply by 2 to both sides, giving us (25+x)/2 * 2= 18 * 2 25+x = 36 subtract by 25 to both sides, 25+x -25 = 36-25 x = 11

OpenStudy (anonymous):

okay, so what do I put, sorry I'mm really confusedd.

OpenStudy (anonymous):

11?

OpenStudy (raden):

oh yeah, ur question just to find x only :) it is x = 11

OpenStudy (anonymous):

okay, thank you , i have one coming now :)